OpenStudy (anonymous):

Write the equation for the line in slope-intercept form from the table below. x 1 2 3 4 5 y 1 -8 -17 -26 -35

OpenStudy (anonymous):

Slope intercept form is y=mx+b

OpenStudy (anonymous):

and if you look at the pattern the outputs are going down by 9: 1-9=8, -8-9=-17 etc. so the slope is -9 or m=-9

OpenStudy (anonymous):

then the y-intercept is the output of 0 to do that you add 9 to the 1 and you find out that the output of 0 is 10 meaning the slope intercept form of your table is: y=-9x+10

OpenStudy (anonymous):

and the table is now: (0,10) (1,1) (2,-8) (3,-17) (4,-26) (5,-35)